# Universal platform of point-gap topological phases from topological materials

### v4

@@ Abstract
Remarkably, the non-Hermiticity enables another extension of topology, the point-gap topology defined by $E\neq E_{\rm P}$ Gong et al. (2018); Kawabata et al. (2019). Typically, a spectrum with nontrivial point-gap topology surrounds the reference point $E_{\rm P}$ in the complex energy plane, and thus the point-gap topology is distinct from any Hermitian-like topology. The topological classification of point-gapped Hamiltonians has been established Zhou and Lee (2019); Kawabata et al. (2019), and the physical consequences of the point-gap topological phases have been explored Gong et al. (2018); D. S. Borgnia, and A. J. Kruchkov, and R.-J. Slager (2020); N. Okuma, and M. Sato (2019); Okuma et al. (2020); Zhang et al. (2020); Okuma and Sato (2021); Kawabata et al. (2021); Yang et al. (2021); Vecsei et al. (2021); Denner et al. (2021); Hu et al. (2022); Nakamura et al. (2022); Denner et al. (2021); Lee et al. (2019a); Terrier and Kunst (2020); Bessho and Sato (2021); Lee et al. (2019b); Luo and Zhang (2019); Li et al. (2020); Okugawa et al. (2020); Kawabata et al. (2020); Fu et al. (2021); Shiozaki and Ono (2021); Palacios et al. (2021); Kim and Park (2021); Zhang et al. (2021); Ghorashi et al. (2021a, b); Zou et al. (2021); Li et al. (2022); Zhu and Gong (2022); Denner and Schindler (2023); Liu and Fulga (2023); Shang et al. (2022); Zhu and Gong (2023); Roccati et al. (2023); Nakai et al. (2023). In particular, it has been shown that the non-Hermitian skin effects originate from one-dimensional point-gap topological numbers, i.e. the spectral winding number Okuma et al. (2020); Zhang et al. (2020) or the $\mathbb{Z}_{2}$ number Okuma et al. (2020). Also, higher-dimensional point-gap topological phases may support non-Hermitian skin modes localized at topological defects, mimicking the anomaly-induced catastrophes Okuma et al. (2020); Okuma and Sato (2021); Kawabata et al. (2021); Denner and Schindler (2023); Okuma and Sato (2023). Depending on the dimension and symmetry of the system, higher-dimensional point-gap topological phases may also host boundary modes Yang et al. (2021); Vecsei et al. (2021); Denner et al. (2021); Hu et al. (2022); Nakamura et al. (2022); Denner and Schindler (2023), and one of such point-gap topological phases is called an exceptional topological insulator Denner et al. (2021). Furthermore, for the fundamental symmetry classes called ${}^{\dagger}$ classes (see below), the correspondence between $d$-dimensional point-gap topological phases and $(d-1)$-dimensional anomalous gapless modes was suggested Lee et al. (2019a), and later proved as the extended Nielsen-Ninomiya theorem Bessho and Sato (2021).
@@
To prove the above statement, we first clarify the symmetry of the system. The fundamental onsite symmetries for topological insulators and superconductors are time-reversal (TRS), particle-hole (PHS), and their combination, chiral symmetry (CS). These Altland-Zirnbauer (AZ) symmetries Altland and Zirnbauer (1997) protect topological gapless boundary modes of topological insulators and superconductors Schnyder et al. (2008); Kitaev (2009); Ryu et al. (2010); Chiu et al. (2016). In the presence of the coupling to the environment, however, we cannot retain these symmetries in their original form. Causal fermionic theories require other onsite symmetries intrinsic to non-Hermitian systems Lieu et al. (2020); Yoshida et al. (2020), which we call ${}^{\dagger}$ symmetries Kawabata et al. (2019). The ${}^{\dagger}$ symmetries consist of ${}^{\dagger}$: $H=U_{T}H^{T}U_{T}^{\dagger}$, $U_{T}U_{T}^{*}=\pm 1$, ${}^{\dagger}$: $H=-U_{C}H^{*}U_{C}^{\dagger}$, $U_{C}U_{C}^{*}=\pm 1$, ${}^{\dagger}$: $H=-U_{\Gamma}H^{\dagger}U_{\Gamma}^{\dagger}$, $U_{\Gamma}U_{\Gamma}=1$, where $H$ is the Hamiltonian, and $U_{T}$, $U_{C}$ and $U_{\Gamma}$ are unitary matrices Kawabata et al. (2019). For a Hermitian $H$, the ${}^{\dagger}$ symmetries coincide with the original AZ symmetries. The presence and absence of the ${}^{\dagger}$ symmetries define the ten-fold ${}^{\dagger}$ classes Kawabata et al. (2019). See Table 1. One can easily check that the onsite decay constant term $-i\Gamma{\bm{1}}$ due to dissipation respects the ${}^{\dagger}$ symmetries  [^4] For details, see Sec.S3 in Supplement Material, which includes Ref. Yoshida et al. (2020)..
A critical mathematical result for our theory is the extended Nielsen-Ninomiya theorem Bessho and Sato (2021), which holds for systems in the ${}^{\dagger}$ classes. The theorem relates the bulk gapless points at the energies $E_{\alpha}$ with the topological charges $\nu_{\alpha}$ to the point-gap topological number $n$ at the reference energy $E_{\rm P}$,
@@
Now we come back to our system, i.e. a $d$-dimensional topological insulator or superconductor with the open boundary condition in the $x_{d}$-direction. Our system belongs to an ${}^{\dagger}$ symmetry class when the system is coupled to the environment at $x_{d}=1$. Regarding the site index in the $x_{d}$-direction as an internal degree of freedom, we can identify the system as a $(d-1)$ dimensional system with ”bulk” gapless states with the internal index $x_{d}=1,L_{d}$. Then, by the coupling to the environment at $x_{d}=1$, the $(d-1)$-dimensional bulk gapless states at $x_{d}=1$ have a different decay constant than those at $x_{d}=L_{d}$. Therefore, from the extended Nielsen-Ninomiya theorem in Eq.(6), there exists a region of $E_{\rm P}$ where the $(d-1)$-dimensional point-gap topological number $n$ becomes non-zero. The non-zero value of $n$ is given by the $d$-dimensional bulk topological number of the original topological insulator/superconductor since the total topological charge of the gapless states at $x_{d}=1$ coincides with it up to sign.
Table 1: Point-gap topological table for insulators and superconductors with dissipation at $x_{d}=1$. Here $\delta=(d-1)-D$, where $d$ is the spatial dimension of the topological insulators and superconductors and $D$ is the dimension of a sphere surrounding a topological defect. The topological defect goes through the $x_{d}$-direction. The superscripts SE and BS indicate the topological numbers predicting non-Hermitian skin effects and boundary states, respectively. See also Ref.Nakamura et al. (2022). For $D=0$, this table reproduces that for $(d-1)$-dimensional point-gap phases in ${}^{\dagger}$ classes in Kawabata et al. (2019).
${}^{\dagger}$ class  ${}^{\dagger}$  ${}^{\dagger}$   CS  $\delta=0$  $\delta=1$  $\delta=2$
@@
Predictions.– The non-zero $(d-1)$-dimensional point-gap topological number $n$ gives rise to several consequences in the physical properties of the system. First, it predicts the appearance of $(d-2)$-dimensional boundary modes or skin modes when imposing the additional open boundary condition on a different direction than $x_{d}$, say the $x_{d-1}$-direction Nakamura et al. (2022). For $d=2$, the non-zero $n$ predicts a second-order non-Hermitian skin effect like the Chern insulator case in Fig.3. The second-order skin modes form a generalized Kramers pair Sato et al. (2012); Kawabata et al. (2019) when the original system has fermionic time-reversal symmetry. For $d=3$, an odd $n$ in class ${}^{\dagger}$ (time-reversal invariant topological superconductor) and a non-trivial $n$ in class ${}^{\dagger}$ (topological insulator) imply the in-gap non-Hermitian skin effects. Still, other non-zero $n$s’ predict boundary modes. (See Table 1 with $D=0$ introduced below.)
Second, the proposed system also may have similar localized modes in the presence of topological defects. The topological defects should go through the $x_{d}$-direction since our theory treats the site index in the $x_{d}$-direction as an internal degree of freedom. Then, we can obtain the point-gap topological table in the presence of such topological defects by generalizing the argument by Teo and Kane Teo and Kane (2010) to point-gap topological phases. See Table 1. In the new table, the space dimension $d-1$ of the point-gap topological phases is replaced by $\delta=(d-1)-D$ $(\geq 0)$, where $D$ $(\leq d-1)$ is the dimension of a sphere surrounding the topological defect. In particular, if we insert the $\pi$-flux ($D=1$) in the $x_{3}$-direction through a three-dimensional time-reversal invariant superconductor with an odd number of the three-dimensional winding number or a three-dimensional time-reversal invariant topological insulator, we have a non-trivial $\mathbb{Z}_{2}^{\rm SE}$ number for class ${}^{\dagger}$ or class ${}^{\dagger}$ with $\delta=1$, in the presence of dissipation at $z=1$. Thus, we obtain the non-Hermitian skin modes localized on the $\pi$-flux.
@@
[…]
@@
S3 S3. ${}^{\dagger}$ symmetry
In this section, we show that AZ symmetry in the total Hamiltonian results in ${}^{\dagger}$ symmetry in the effective Hamiltonian in Sec. S2.
@@
which leads to the transpose version of TRS, which we call ${}^{\dagger}$, in the effective Hamiltonian for the topological material.
@@
respectively, in the effective Hamiltonian. Therefore, AZ symmetry of the total Hamiltonian results in ${}^{\dagger}$ symmetry (${}^{\dagger}$, ${}^{\dagger}$ and CS ) in the effective Hamiltonian Yoshida et al. (2020).

### v5

@@ Abstract
Remarkably, the non-Hermiticity enables another extension of topology, the point-gap topology defined by $E\neq E_{\rm P}$ Gong et al. (2018); Kawabata et al. (2019). Typically, a spectrum with nontrivial point-gap topology surrounds the reference point $E_{\rm P}$ in the complex energy plane, and thus the point-gap topology is distinct from any Hermitian-like topology. The topological classification of point-gapped Hamiltonians has been established Zhou and Lee (2019); Kawabata et al. (2019), and the physical consequences of the point-gap topological phases have been explored Gong et al. (2018); D. S. Borgnia, and A. J. Kruchkov, and R.-J. Slager (2020); N. Okuma, and M. Sato (2019); Okuma et al. (2020); Zhang et al. (2020); Okuma and Sato (2021); Kawabata et al. (2021); Yang et al. (2021); Vecsei et al. (2021); Denner et al. (2021); Hu et al. (2022); Nakamura et al. (2022); Denner et al. (2021); Lee et al. (2019a); Terrier and Kunst (2020); Bessho and Sato (2021); Lee et al. (2019b); Luo and Zhang (2019); Li et al. (2020); Okugawa et al. (2020); Kawabata et al. (2020); Fu et al. (2021); Shiozaki and Ono (2021); Palacios et al. (2021); Kim and Park (2021); Zhang et al. (2021); Ghorashi et al. (2021a, b); Zou et al. (2021); Franca et al. (2022); Li et al. (2022); Zhu and Gong (2022); Denner and Schindler (2023); Liu and Fulga (2023); Shang et al. (2022); Zhu and Gong (2023); Roccati et al. (2023); Nakai et al. (2023). In particular, it has been shown that the non-Hermitian skin effects originate from one-dimensional point-gap topological numbers, i.e. the spectral winding number Okuma et al. (2020); Zhang et al. (2020) or the $\mathbb{Z}_{2}$ number Okuma et al. (2020). Also, higher-dimensional point-gap topological phases may support non-Hermitian skin modes localized at topological defects, mimicking the anomaly-induced catastrophes Okuma et al. (2020); Okuma and Sato (2021); Kawabata et al. (2021); Denner and Schindler (2023); Okuma and Sato (2023). Depending on the dimension and symmetry of the system, higher-dimensional point-gap topological phases may also host boundary modes Yang et al. (2021); Vecsei et al. (2021); Denner et al. (2021); Hu et al. (2022); Nakamura et al. (2022); Denner and Schindler (2023), and one of such point-gap topological phases is called an exceptional topological insulator Denner et al. (2021). Furthermore, for the fundamental symmetry classes called AZ classes (see below), the correspondence between $d$-dimensional point-gap topological phases and $(d-1)$-dimensional anomalous gapless modes was suggested Lee et al. (2019a), and later proved as the extended Nielsen-Ninomiya theorem Bessho and Sato (2021).
@@
To prove the above statement, we first clarify the symmetry of the system. The fundamental onsite symmetries for topological insulators and superconductors are time-reversal (TRS), particle-hole (PHS), and their combination, chiral symmetry (CS). These Altland-Zirnbauer (AZ) symmetries Altland and Zirnbauer (1997) protect topological gapless boundary modes of topological insulators and superconductors Schnyder et al. (2008); Kitaev (2009); Ryu et al. (2010); Chiu et al. (2016). In the presence of the coupling to the environment, however, we cannot retain these symmetries in their original form. Causal fermionic theories require other onsite symmetries intrinsic to non-Hermitian systems Lieu et al. (2020); Yoshida et al. (2020), which we call AZ symmetries Kawabata et al. (2019). The AZ symmetries consist of TRS: $H=U_{T}H^{T}U_{T}^{\dagger}$, $U_{T}U_{T}^{*}=\pm 1$, PHS: $H=-U_{C}H^{*}U_{C}^{\dagger}$, $U_{C}U_{C}^{*}=\pm 1$, CS: $H=-U_{\Gamma}H^{\dagger}U_{\Gamma}^{\dagger}$, $U_{\Gamma}U_{\Gamma}=1$, where $H$ is the Hamiltonian, and $U_{T}$, $U_{C}$ and $U_{\Gamma}$ are unitary matrices Kawabata et al. (2019). For a Hermitian $H$, the AZ symmetries coincide with the original AZ symmetries. The presence and absence of the AZ symmetries define the ten-fold AZ classes Kawabata et al. (2019). See Table 1. One can easily check that the onsite decay constant term $-i\Gamma{\bm{1}}$ due to dissipation respects the AZ symmetries  [^4] For details, see Sec.S3 in Supplement Material, which includes Ref. Yoshida et al. (2020)..
A critical mathematical result for our theory is the extended Nielsen-Ninomiya theorem Bessho and Sato (2021), which holds for systems in the AZ classes. The theorem relates the bulk gapless points at the energies $E_{\alpha}$ with the topological charges $\nu_{\alpha}$ to the point-gap topological number $n$ at the reference energy $E_{\rm P}$,
@@
Now we come back to our system, i.e. a $d$-dimensional topological insulator or superconductor with the open boundary condition in the $x_{d}$-direction. Our system belongs to an AZ symmetry class when the system is coupled to the environment at $x_{d}=1$. Regarding the site index in the $x_{d}$-direction as an internal degree of freedom, we can identify the system as a $(d-1)$ dimensional system with ”bulk” gapless states with the internal index $x_{d}=1,L_{d}$. Then, by the coupling to the environment at $x_{d}=1$, the $(d-1)$-dimensional bulk gapless states at $x_{d}=1$ have a different decay constant than those at $x_{d}=L_{d}$. Therefore, from the extended Nielsen-Ninomiya theorem in Eq.(6), there exists a region of $E_{\rm P}$ where the $(d-1)$-dimensional point-gap topological number $n$ becomes non-zero. The non-zero value of $n$ is given by the $d$-dimensional bulk topological number of the original topological insulator/superconductor since the total topological charge of the gapless states at $x_{d}=1$ coincides with it up to sign.
Table 1: Point-gap topological table for insulators and superconductors with dissipation at $x_{d}=1$. Here $\delta=(d-1)-D$, where $d$ is the spatial dimension of the topological insulators and superconductors and $D$ is the dimension of a sphere surrounding a topological defect. The topological defect goes through the $x_{d}$-direction. The superscripts SE and BS indicate the topological numbers predicting non-Hermitian skin effects and boundary states, respectively. See also Ref.Nakamura et al. (2022). For $D=0$, this table reproduces that for $(d-1)$-dimensional point-gap phases in AZ classes in Kawabata et al. (2019).
AZ class  TRS  PHS   CS  $\delta=0$  $\delta=1$  $\delta=2$
@@
Predictions.– The non-zero $(d-1)$-dimensional point-gap topological number $n$ gives rise to several consequences in the physical properties of the system. First, it predicts the appearance of $(d-2)$-dimensional boundary modes or skin modes when imposing the additional open boundary condition on a different direction than $x_{d}$, say the $x_{d-1}$-direction Nakamura et al. (2022). For $d=2$, the non-zero $n$ predicts a second-order non-Hermitian skin effect like the Chern insulator case in Fig.3. The second-order skin modes form a generalized Kramers pair Sato et al. (2012); Kawabata et al. (2019) when the original system has fermionic time-reversal symmetry. For $d=3$, an odd $n$ in class DIII (time-reversal invariant topological superconductor) and a non-trivial $n$ in class AII (topological insulator) imply the in-gap non-Hermitian skin effects. Still, other non-zero $n$s’ predict boundary modes. (See Table 1 with $D=0$ introduced below.)
Second, the proposed system also may have similar localized modes in the presence of topological defects. The topological defects should go through the $x_{d}$-direction since our theory treats the site index in the $x_{d}$-direction as an internal degree of freedom. Then, we can obtain the point-gap topological table in the presence of such topological defects by generalizing the argument by Teo and Kane Teo and Kane (2010) to point-gap topological phases. See Table 1. In the new table, the space dimension $d-1$ of the point-gap topological phases is replaced by $\delta=(d-1)-D$ $(\geq 0)$, where $D$ $(\leq d-1)$ is the dimension of a sphere surrounding the topological defect. In particular, if we insert the $\pi$-flux ($D=1$) in the $x_{3}$-direction through a three-dimensional time-reversal invariant superconductor with an odd number of the three-dimensional winding number or a three-dimensional time-reversal invariant topological insulator, we have a non-trivial $\mathbb{Z}_{2}^{\rm SE}$ number for class DIII or class AII with $\delta=1$, in the presence of dissipation at $z=1$. Thus, we obtain the non-Hermitian skin modes localized on the $\pi$-flux.
@@
Our scheme applies to any topological materials in the original topological periodic table Schnyder et al. (2008); Kitaev (2009); Ryu et al. (2010); Chiu et al. (2016). For instance, by connecting a metal to an edge of a quantum Hall state in graphene, we can realize a point-gap topological phase similar to Fig.3. The resulting non-Hermitian skin effect can be observed as the chiral tunneling effect Yi and Yang (2020). Another candidate is a topological superconducting nanowire with a Zeeman field Sato and Ando (2017). By coupling a lead to one of the ends of the nanowire, the system displays a point-gap topological phase in 0D class D. We can also use the topological insulator Bi2Se3 and the variants Ando (2013) to similarly realize a point-gap topological phase in 2D class AII.
@@
S3 S3. AZ symmetry
In this section, we show that AZ symmetry in the total Hamiltonian results in AZ symmetry in the effective Hamiltonian in Sec. S2.
@@
which leads to the transpose version of TRS, which we call TRS, in the effective Hamiltonian for the topological material.
@@
respectively, in the effective Hamiltonian. Therefore, AZ symmetry of the total Hamiltonian results in AZ symmetry (TRS, PHS and CS ) in the effective Hamiltonian Yoshida et al. (2020).